# Dark Sector Tunneling Field Potentials for a Dark Big Bang

### v1

@@ IV Results
For our analysis, we have fixed $\lambda$ = 1 to allow for the decay of $\phi$ to be very efficient (this choice was made for each of the benchmark Dark Big Bangs considered in [1]). Future research could look into how the parameter space changes with $\lambda$. For reasons outlined in Appendix A, we will discuss the parameter space in two regions. Region 1, which we have dubbed “the sliver”, rides the singularity in $S$ (where $\mu^{2}=2\lambda m^{2}$). This region’s proximity to the singularity makes the PT parameters very sensitive to choices of $m$ and $\mu$. Furthermore, this parameter space is so small that it cannot be visualized over large ranges of $m$ and $\mu$. A second region of parameter space (Region 2) opens up when $\mu\gg$ $m$ (see Appendix A). Region 1 of the available parameter space for a DBB is shown in Figs. 3 and 4. Both regions are shown in Figs. 5 and 7.
Throughout much of the available parameter space, the bound on the latest a DBB can occur (Lyman-$\alpha$ bound) is usually much weaker than the upper and lower bounds on the strength of the DBB $\alpha$. However, this bound does play a role at small $m$ and $\mu$ as shown in Fig. 5 and can be visually inferred by the gentler slope of the green region in Fig. 3. The Lyman-$\alpha$ bound can be turned into a constraint between $\alpha$ and $\Delta$V by combining Eqs. 7 and 13: $\alpha\lesssim 4.24\times 10^{12}\Delta V$ MeV-4. Using this constraint and Eq. 11, the parameter space is ruled out towards small $m$ and $\mu$ values when
$\Delta V(m,\mu)\approx\frac{1.37\times 10^{-19}}{T_{*}(m,\mu)}\rm{MeV}^{4}=\alpha(m,\mu)$ (20)
which can be solved for $m$ and $\mu$ numerically. The values of $m$ and $\mu$ for which the parameter space is ruled out by the temperature bound are shown in Tab. 1.
@@
Figure 3: The available parameter space for a DBB when $\mu\gtrsim m$. The bounds have been recast from Fig. 2 to $m$-$\mu$ space using Eqs. 18 and 19. Also plotted is the constraint that $\mu>\sqrt{2\lambda m}$ with $\lambda$=1. The white region shows tunneling field parameter choices for a DBB consistent with the bounds and widens (narrows) beyond the figure to the right (left). While it appears that the parameter space is closed for most of the figure, Fig. 4 will show that the parameter space remains open throughout this range and is too narrow to visualize at this scale.
@@
Figure 4: Magnified regions of the sliver parameter space shown in Fig. 3. The regions chosen correspond to the four benchmark Dark Big Bangs considered in [1]. Interestingly, we found that the choices of parameters for BP1 (Panel 1) and Dark-Zillas (Panel 4) violate the upper bound on $\alpha$. We believe this is a result of compounding rounding errors due to sensitivity of this region. These discrepancies do not significantly impact the results of [1], as the parameters can be adjusted slightly to produce the same PT characteristics used in their analysis.
@@
Figure 5: Region 2 of the available parameter space is shown in white with the relevant bounds as in Fig. 2. Using a log scale on both axes, the sliver region is imperceptible but remains open and is represented by the gray boundary line. Above $m\approx 6.1\times 10^{9}$, the lower (orange) bound on $\alpha$ plays no role and the parameter space is bounded on both sides by the upper (blue) bound.
The two separated regions of parameter space are a consequence of the minimum of $\alpha$ at constant $m$ dropping below the lower bound before returning back through the parameter space. As $m$ and $\mu$ increase, the minimum value of $\alpha$ increases. Eventually, the minimum of $\alpha$ crosses above the lower bound and the two parameter spaces merge. Above this point, the parameter space is bounded on both sides by the upper bound on $\alpha$ until not even the minimum value of $\alpha$ crosses below the upper bound. The critical points are summarized in Tab. 1.
@@ V Discussion
We are particularly interested in looking at the qualitative differences between the two regions. In Region 2, $\mu\gg m$, making analytic analysis much simpler than in the sliver. Additionally, there has yet to be a benchmark DBB considered with input parameters drawn from this region. Given these considerations, this section will first look at the features of tunneling potentials from Region 2 and then consider benchmark DBBs using input parameters from this region.
@@ V.1 Tunneling Potentials
Looking at the quartic potential in Eq. 2, one can deduce the effects $m,\mu$, and $\lambda$ have on the shape of the potential. As $\phi$ increases from zero, $V(\phi)$ first increases $\sim m^{2}\phi^{2}$, then decreases $\sim\mu\phi^{3}$, and finally increases $\sim\lambda\phi^{4}$. So throughout Region 2, where $\mu\gg m$, the potential barrier protecting the stability of the false vacuum is the smallest feature of the tunneling potential, as shown in Fig. 6. This section will show that although $V_{\rm{b}}$ is small throughout Region 2, the tunneling potentials from this region can successfully provide the initial conditions for a first-order phase transition in the dark sector.
Figure 6: Benchmark tunneling field potential from Region 2 of the parameter space. The potential shown here corresponds to the input parameters of DW 1 shown in Tab. 2. Although the potential barrier is imperceptible, it is large enough to withstand thermal fluctuations. For this potential, $V_{\rm{b}}\sim 10^{-77}$ MeV4 while $T_{\rm{GH}}^{4}\sim 10^{-106}$ MeV4.
Additional constraints on the tunneling field potential come from the stability of the dark false vacuum. While the tunneling field is trapped in the metastable minimum of the potential, the dark sector goes through a period of inflation. For the dark false vacuum to remain stable during inflation, fluctuations caused by the Gibbons-Hawking temperature $T_{\rm{GH}}$ must not exceed the potential barrier $(T_{\rm{GH}})^{4}\lesssim V_{\rm{b}}$ [1, 24]. The Gibbons-Hawking temperature is proportional to the Hubble parameter during inflation $T_{\rm{GH}}=H_{I}/2\pi$ [25]. We then have the condition [^3] A weaker constraint on the stability of the false vacuum $\Delta V\lesssim 3M_{p}^{2}T_{*}^{2}$ is considered in [1].
$V_{\rm{b}}\gtrsim\left(\frac{1}{2\pi}\right)^{4}\left(\frac{\Delta V}{3M_{p}^{2}}\right)^{2}$ (21)
where the Hubble parameter of the dark sector during inflation is defined with $\rho_{\rm{DS}}=\Delta V$. In Region 2, we can safely use analytic approximations for $V_{\rm{b}}$ and $\Delta V$. It can be shown that for $\mu\gg\rm{m}$,
@@
$V_{\rm{b}}\approx\frac{576m^{6}}{31104\mu^{2}}$ (23)
Eqs. 22 and 23 show that $\Delta V>V_{\rm{b}}$ throughout Region 2, ensuring the tunneling field will indeed decay and not remain in the false vacuum forever. Yet it is important to check that the barrier height is large enough to remain stable during inflation because of the $V_{\rm{b}}\sim\mu^{-2}$ dependence. Using these approximations, we find that all of Region 2 produces tunneling field potentials with barrier heights large enough to avoid destabilization from fluctuations caused by $T_{\rm{GH}}$, i.e. the Gibbons-Hawking bound is weaker than the upper limit on $\alpha$, as shown in Fig. 7.
Because $m\ll\mu$ throughout Region 2, there appears to be risk of isocurvature perturbations [26] being generated in the dark matter content (see [1] for details). We find that throughout Region 2, $m>H_{I}$ where as before $H_{I}=\sqrt{\frac{\Delta V}{3M_{\rm{p}}^{2}}}$ and $m$ is the mass of the tunneling field in the false vacuum. Therefore isocurvature perturbations are not induced since quantum fluctuations in $\phi$ are heavily suppressed. Fig. 7 shows that imposing the bounds for false vacuum stability (Gibbons-Hawking) and the upper limit on the strength of the DBB (CMB $\Delta N_{\rm{eff}}$) ensure that isocurvature perturbations are safely avoided throughout Region 2.
@@
Figure 7: The available parameter space for a DBB consistent with observations, including constraints on the scale of (dark sector) inflation. The open parameter space is the same as Fig. 5 with an additional bound from the Trans-Planckian Censorship Conjecture (TCC), since the bounds from unwanted isocurvature perturbations and false vacuum stability considerations are both weaker than the upper bound on $\alpha$. The TCC bound closes the parameter space above it entirely at the ($m,\mu$) value shown in Tab. 1.
Appendix B shows calculated PT and tunneling field parameters for the critical points in Tab. 1. A notable result of these calculations is that a DBB as late as $\approx$ 32.5 days (the latest a DBB can occur) after the Hot Big Bang can be realized in both Region 1 and Region 2. The only significant difference (for a very late DBB) between these two regions is the mass shift $m$ undergoes in transitioning from the false vacuum to the true vacuum: $\mathcal{O}(1)$ in Region 1, $\mathcal{O}(10^{12})$ in Region 2.
@@ V.2 Benchmark Dark Big Bangs
$m_{\chi}\lesssim\gamma_{w}m_{\phi}\simeq 2\frac{\Delta V}{m(\Delta\phi)^{2}}\frac{M_{\rm{p}}}{\sqrt{g_{\rm{eff}}(T_{*})T_{*}^{4}}}m_{\phi}$ (25)
@@
The largest deviations from the freeze out of standard WIMPs come from the dark and visible sectors having different temperatures. In the dark WIMP case, the equilibrium abundance $Y_{\chi,\rm{eq}}$ depends on $T_{\rm{DS}}$, as does when $\chi$ is considered to be relativistic $m_{\chi}/T_{\rm{DS}}>1$. Since entropy is conserved in both sectors and, for the dark WIMP case, the dark sector always has relativistic degrees of freedom, the ratio of the temperatures of the two sectors is constant up to changes in the degrees of freedom [1]
@@
Eq. 28 can be used to put $T_{\rm{DS}}$ in terms of $T$. For the benchmark cases considered here, we took $g_{\rm{DS},*}=2$ if $\chi$ is produced relativistically and $g_{\rm{DS}}=1$ once $\chi$ becomes non-relativistic ($\xi$ will always contribute to the dark sector degrees of freedom). In our calculations, we approximated $g_{\rm{DS}}$ with a continuous step function of the form
@@
where $g_{\rm{DS},i}$ $(g_{\rm{DS},f})$ is the initial (final) relativistic degrees of freedom in the dark sector, $T_{\rm{nr}}$ is the visible sector temperature at which $\chi$ becomes non relativistic ($m_{\chi}/T_{\rm{DS}}=1$) [^4] We approximated this temperature by using Eq. 28 when all of $\chi$ is non relativistic ($g_{\rm{DS}}(T_{\rm{DS}})=1$)., and K is a constant determining how fast $\chi$ becomes non-relativistic [^5] An appropriate choice has 80% of $\chi$ become non-relativistic by $T_{\rm{DS}}\approx\frac{1}{6}m_{\chi}$ [17].. The benchmark cases considered are shown in Tab. 2.
@@
$\langle\sigma_{\chi\chi\rightarrow\xi\xi}v\rangle$ [$\rm{cm^{3}s^{-1}}$] $1.3\times 10^{-26}$ $1.3\times 10^{-27}$
@@
Table 2: Benchmark Dark Big Bangs producing Dark WIMP (DW) dark matter with input parameters from Region 2. The input parameters were chosen to feature phase transitions that occur at different times in the universe, are strong enough to fall within the sensitivity bounds of upcoming GW surveys, and produce the correct relic abundance assuming Dark WIMPs make up all the dark matter in the universe. Benchmark cases featuring the same Dark WIMP masses with parameters chosen from Region 1 can be found in Ref. [1].
The evolution of $\rho_{\chi}$ for each benchmark case is shown in Fig. 8. Once $Y_{\chi}$ is solved for numerically, the dark WIMP energy density is given by $\rho_{\chi}=m_{\chi}Y_{\chi}s$ in the non relativistic regime. The dark WIMP energy density evolution while the particles are in thermal equilibrium and (for the most part) relativistic is
@@
where $K_{1}$ is a modified Bessel function of the second kind and $s$ is the entropy density of the visible sector $s=\frac{2\pi^{2}}{45}g_{\rm{s}}(T)T^{3}$. Eq. 30 is found by numerically integrating the Maxwell-Boltzmann phase space distributions for $n_{\chi,\rm{eq}}$ and $\rho_{\chi,\rm{eq}}$ [^6] As in Ref. [1], we approximate the Bose-Einstein distribution with a Maxwell-Boltzmann distribution..
Figure 8: Evolution of the Dark WIMP energy density for the two benchmark cases shown in Tab. 2. In both cases, the dark matter is produced relativistically and redshifts like $\rho_{\chi}\propto T^{4}$. The Dark WIMPs freeze out as cold relics and join the standard $\Lambda$CDM evolution track, redshifting like $\rho_{\chi}\propto T^{3}$ and producing the correct relic abundance of dark matter in the late universe.
@@
at the time of the DBB. The energy density of gravity waves will redshift like $\rho_{\rm{GW}}\sim a^{-4}$ [1]. The projected sensitivity windows of the International Pulsar Timing Array (IPTA) [31] and the Square Kilometer Array (SKA) [32] make strong and late DBB’s favorable for detection [1]. Both benchmark Dark Big Bangs fall withing IPTA and SKA sensitivity windows.
VI Conclusion

### v3

@@ IV Results
For our analysis, we have fixed $\lambda$ = 1 to allow for the decay of $\phi$ to be very efficient (this choice was made for each of the benchmark Dark Big Bangs considered in [1]). Future research could look into how the parameter space changes with $\lambda$. For reasons outlined in Appendix A, we will discuss the parameter space in two regions. Region 1, which we have dubbed “the sliver”, rides the singularity in $S$ (where $\mu^{2}=2\lambda m^{2}$). This region’s proximity to the singularity makes the phase transition (PT) parameters very sensitive to choices of $m$ and $\mu$. Furthermore, this region of parameter space is so narrow that it cannot be visualized over large ranges of $m$ and $\mu$. We find that a second region of parameter space (Region 2) opens up when $\mu\gg$ $m$ (see Appendix A). Region 1 of the available parameter space for a DBB is shown in Figs. 3 and 4. Both regions are shown in Figs. 5 and 8.
@@
Figure 3: The available parameter space for a DBB when $\mu\gtrsim m$ (Region 1). The bounds have been recast from Fig. 2 to $m$-$\mu$ space using Eqs. 18 and 19. Also plotted is the constraint that $\mu>\sqrt{2\lambda m}$ with $\lambda=1$. The white region shows tunneling field parameter choices for a DBB consistent with $\Lambda$CDM and widens (narrows) beyond the figure to the right (left). While it appears that the parameter space is closed for most of the figure, Fig. 4 will show that the parameter space remains open throughout this range and is too narrow to visualize at this scale.
Throughout much of the parameter space, the bound on the latest a DBB can occur (Lyman-$\alpha$ bound, e.g. green swath) is weaker than the upper (CMB $\Delta N_{\rm{eff}}$, e.g. blue excluded region) and lower (Dark Matter, e.g. orange excluded region) bounds on the strength of the DBB $\alpha$. However, this bound does play a role at small $m$ and $\mu$ as shown in Fig. 5. The Lyman-$\alpha$ bound can be turned into a constraint between $\alpha$ and $\Delta$V by combining Eqs. 7 and 13: $\alpha\lesssim 4.24\times 10^{12}\Delta V$ MeV-4, where we used $g_{\rm{eff}}(6.8\times 10^{-4}$ MeV) $=3.35$. Using this constraint and Eq. 11, the parameter space is ruled out towards small $m$ and $\mu$ values when the Lyman-$\alpha$ constraint is equal to the Dark Matter bound
$4.24\times 10^{12}\left(\frac{\Delta V(m,\mu)}{\rm{MeV}^{4}}\right)\simeq 5.8\times 10^{-7}\left(\frac{\rm{MeV}}{T_{*}(m,\mu)}\right)=\alpha(m,\mu)$ (20)
which can be solved for $m$ and $\mu$ numerically (we have set $\lambda=1$). The values of $m$ and $\mu$ below which the parameter space is ruled out by the temperature bound are shown in the top two rows of Tab. 1.
@@
Figure 4: Four different magnifications into Region I of the parameter space, which was partially shown Fig. 3. The regions chosen correspond to the four benchmark Dark Big Bangs considered in [1]. Interestingly, we found that the choices of parameters for BP1 (Panel 1) and Dark-Zillas (Panel 4) slightly violate the upper bound on $\alpha$. These discrepancies do not significantly impact the results of [1], as the parameters can be adjusted slightly to produce the same phase transition characteristics used in their analysis.
@@
Figure 5: Full parameter space for tunneling potential parameters that lead to a DBB consistent with observations. At low $m$ and $\mu$ values, Region 1 and and Region 2 are separated by the lower bound on $\alpha$ (i.e. orange excluded region). On this log scale, Region 1 is barely visible, yet remains open and is represented by the white line just above the $\mu^{2}>2m^{2}$ region. For $m\gtrsim 6.1\times 10^{9}$, the lower (orange) bound on $\alpha$ plays no role and the parameter space is bounded on both sides by the upper (blue) bound.
In Fig. 5 we identify the part of the ($\mu$ vs. $m$) parameter space that leads to a DBB which is consistent with all experimental bounds. Note that the white region (allowed) can be naturally split into two. First, the very narrow sliver (Region 1) riding along the $\mu^{2}=2m^{2}$ line, and open from $\mu\sim 0.001$ to $\mu\sim 10^{10}$. This region was previously explored in [1]. Note a second, much wider allowed region (Region 2). The existence of Region 2, which was not explored in [1], is due to the value of $\alpha$ at constant $m$ increasing back through the bounds as $\mu$ increases (see Appendix A for details). This behavior can be seen in Fig. 6, which shows plots of $\alpha$ and $T_{*}g_{\rm{eff}}^{1/4}$ (Eq. 19 and Eq. 18) for fixed $m$ (taking vertical slices of Figs. 345). We define Region 2 as the available parameter space when $\mu\gg m$ (corresponding to the portions of the plots in Fig. 6 with gentler slopes). As both $m$ and $\mu$ increase, the minimum value of $\alpha$ increases, eventually evading the lower bound ($m_{3}$ in the left panel of Fig. 6). For $m$ and $\mu$ greater than this point (listed in row 3 in Tab. 1), the parameter space is bounded on both sides by the upper bound on $\alpha$ (blue region). The parameter space closes when the minimum value of $\alpha$ is larger than the upper bound. The critical points, defining where regions of the parameter space open or close, are summarized in Tab. 1.
In comparing Figs. 3 and 4 with Fig. 5, there is a change in the hierarchy of the bounds on a DBB. Namely, for Region 1 the CMB $\Delta N_{\rm{eff}}$ bound (blue) rules out larger ($m,\mu$) values than the Lyman-$\alpha$ bound (green), whereas for Region 2, the situation is reversed for most of the parameter space. The hierarchy flip is explained by Fig. 6. For each $m$ value plotted in Fig. 6, when $\mu\approx m$, $\alpha$ violates the CMB $\Delta N_{\rm{eff}}$ bound and $T_{*}g_{\rm{eff}}^{1/4}$ violates the Lyman-$\alpha$ bound. Increasing $\mu$, $T_{*}g_{\rm{eff}}^{1/4}$ increases into the allowed region, quickly followed by $\alpha$ decreasing into the allowed region. This behavior is shown in Figs. 3 and 4 where the Lyman-$\alpha$ bound is weaker than the CMB $\Delta N_{\rm{eff}}$ bound throughout Region 1. In Fig. 5, the corresponding section of parameter space follows the $\mu^{2}>2m^{2}$ boundary (the sliver). Looking closely, the CMB $\Delta N_{\rm{eff}}$ bound can be seen (the blue line along the gray shaded region). We have chosen not to plot the Lyman-$\alpha$ bound along the gray region (as it is shown in Figs. 3 and 4) because it is weaker than the blue CMB $\Delta N_{\rm{eff}}$ bound and obscures the edge of the parameter space. Going back to Fig. 6, as $\mu$ continues to increase, $\alpha$ returns to violate the CMB $\Delta N_{\rm{eff}}$ bound long before $T_{*}g_{\rm{eff}}^{1/4}$ returns to violate the Lyman-$\alpha$ bound. Hence the hierarchy flip of the bounds as we move vertically from below Region 1 to above Region 2.
Figure 6: Left Panel: Values of $\alpha$ for fixed $m$. As before, $\alpha$ is bounded above by the CMB $\Delta N_{\rm{eff}}$ (blue) bound and below by the Dark Matter (orange) bounds. For the three $m$ values shown, the CMB $\Delta N_{\rm{eff}}$ bound is the same and constant. The Dark Matter bound changes for each fixed $m$ value. All values of $\alpha$ for the $m_{3}$ plot are above the lower bound. Right Panel: Temperature of the visible sector at the time of the DBB as a function of $\mu$. We have included the effective relativistic degrees of freedom $g_{\rm{eff}}(T)$ in our definition of temperature as in Sec. III.2. Dark Big Bangs that occur at a temperature below the Lyman-$\alpha$ bound are excluded by the current understanding of structure formation.
@@ V Discussion
We are particularly interested in looking at the qualitative differences between the two regions. In Region 2, defined by $\mu\gg m$, analytic analysis is much simpler than in the sliver (Region 1). Additionally, there has yet to be a benchmark DBB considered with input parameters drawn from this region. Given these considerations, this section will first look at the features of tunneling potentials from Region 2 and then consider benchmark DBBs using input parameters from this region.
@@ V.1 Tunneling Potentials
Looking at the quartic potential in Eq. 2, one can deduce the effects $m,\mu$, and $\lambda$ have on the shape of the potential. As $\phi$ increases from zero, $V(\phi)$ first increases $\sim m^{2}\phi^{2}$, then decreases $\sim\mu\phi^{3}$, and finally increases $\sim\lambda\phi^{4}$. So throughout Region 2, where $\mu\gg m$, the potential barrier protecting the stability of the false vacuum is the smallest feature of the tunneling potential, as shown in Fig. 7. This section will show that although $V_{b}$ is small throughout Region 2, the tunneling potentials from this region can successfully provide the initial conditions for a first-order phase transition in the dark sector.
Figure 7: Benchmark tunneling field potential from Region 2 of the parameter space. The potential shown here corresponds to the input parameters of DW 1 shown in Tab. 2. Although the potential barrier is imperceptible, it is large enough to withstand thermal fluctuations. For this potential, $V_{b}\sim 10^{-77}$ MeV4 while $T_{\rm{GH}}^{4}\sim 10^{-106}$ MeV4.
Additional constraints on the tunneling field potential come from the stability of the dark false vacuum. While the tunneling field is trapped in the metastable minimum of the potential, the dark sector goes through a period of inflation. For the dark false vacuum to remain stable during inflation, the potential barrier ($V_{b}$) must exceed fluctuations caused by the Gibbons-Hawking temperature $T_{\rm{GH}}$ [1, 27]
$V_{b}\gtrsim(T_{\rm{GH}})^{4}.$ (21)
The Gibbons-Hawking temperature is proportional to the Hubble parameter during inflation $T_{\rm{GH}}=H_{I}/2\pi$ [28]. We then have the condition [^3] A weaker constraint on the stability of the false vacuum $\Delta V\lesssim 3M_{p}^{2}T_{*}^{2}$ is considered in [1].
$V_{b}\gtrsim\left(\frac{1}{2\pi}\right)^{4}\left(\frac{\Delta V}{3M_{P}^{2}}\right)^{2}$ (22)
where the Hubble parameter of the dark sector during inflation is defined with $\rho_{\rm{DS}}=\Delta V$. In Region 2, we can safely use analytic approximations for $V_{b}$ and $\Delta V$. It can be shown that for $\mu\gg\rm{m}$
@@
$V_{b}\approx\frac{576m^{6}}{31104\mu^{2}}$ (24)
Eqs. 23 and 24 show that $\Delta V>V_{b}$ throughout Region 2, ensuring the tunneling field will indeed decay and not remain in the false vacuum forever. At this stage we will check that the barrier height is large enough to remain stable during inflation. Using these approximations, we find that all of Region 2 produces tunneling field potentials with barrier heights large enough to avoid destabilization from fluctuations caused by $T_{\rm{GH}}$, i.e. the Gibbons-Hawking bound is weaker than the upper limit on $\alpha$, as shown in Fig. 8.
Because $\mu\gg m$ throughout Region 2, there appears to be risk of isocurvature perturbations [29] being generated in the dark matter content (see [1] for details). We find that throughout Region 2, $m>H_{I}$ where as before $H_{I}=\sqrt{\frac{\Delta V}{3M_{\rm{p}}^{2}}}$ and $m$ is the mass of the tunneling field in the false vacuum. Therefore isocurvature perturbations are not induced since quantum fluctuations in $\phi$ are heavily suppressed [1]. Fig. 8 shows that imposing the bounds for false vacuum stability (Gibbons-Hawking) and the upper limit on the strength of the DBB (CMB $\Delta N_{\rm{eff}}$) ensure that isocurvature perturbations are safely avoided throughout Region 2.
@@
Figure 8: The available parameter space for a DBB consistent with observations, including constraints on the scale of (dark sector) inflation. The open parameter space is the same as Fig. 5 with an additional bound from the Trans-Planckian Censorship Conjecture (TCC), since the bounds from unwanted isocurvature perturbations (yellow region: $m<H_{I}$) and false vacuum stability considerations are both weaker than the upper bound on $\alpha$. The TCC bound closes the parameter space above it entirely at the ($m,\mu$) value shown in Tab. 1.
Appendix B shows calculated PT and tunneling field parameters for the critical points in Tab. 1. A notable result of these calculations is that a DBB as late as $\approx$ 32.5 days (the latest a DBB can occur) after the Hot Big Bang can be realized in both Region 1 and Region 2. The only significant difference between these two regions (for a very late DBB) is the mass shift $m$ undergoes in transitioning from the false vacuum to the true vacuum. In both cases, the mass in the true vacuum $m_{\phi}$ is $\mathcal{O}(10^{-4}$ MeV) yet the starting masses (in the false vacuum) are very different. In Region 2, $m$ in the false vacuum is $\mathcal{O}(10^{-16}$ MeV) while in Region 1, $m$ is $\mathcal{O}(10^{-4}$ MeV) (see Tab. 3 in Appendix B).
@@ V.2 Benchmark Dark Big Bangs
$m_{\chi}\lesssim\gamma_{w}m_{\phi}\simeq 2\frac{\Delta V}{m(\Delta\phi)^{2}}\frac{M_{P}}{\sqrt{g_{\rm{eff}}(T_{*})T_{*}^{4}}}m_{\phi}$ (26)
@@
The largest deviations from the freeze out of standard WIMPs come from the dark and visible sectors having different temperatures. In the Dark WIMP case, the equilibrium abundance $Y_{\chi,\rm{eq}}$ depends on $T_{\rm{DS}}$, as does when $\chi$ is considered to be relativistic $m_{\chi}/T_{\rm{DS}}<1$. Since entropy is conserved in both sectors and, for the dark WIMP case, the dark sector always has relativistic degrees of freedom, the ratio of the temperatures of the two sectors is constant up to changes in the degrees of freedom [1]
@@
Eq. 29 is used to put $T_{\rm{DS}}$ in terms of $T$. For the benchmark cases considered here, we took $g_{\rm{DS},*}=2$ if $\chi$ is produced relativistically and $g_{\rm{DS}}=1$ once $\chi$ becomes non-relativistic ($\xi$ will always contribute to the dark sector degrees of freedom).
To incorporate $g_{\rm{DS}}(T_{\rm{DS}})$ into our calculations, we created a continuous step function modeled after the behavior of $g_{\rm{eff}}$(T) in the visible sector. The step function goes from an initial value of the relativistic degrees of freedom in the dark sector $g_{\rm{DS},i}$ to a final value $g_{\rm{DS},f}$ with the transition centered at $T_{\rm{nr}}$. The steepness of the transition is determined by the parameter $K$, with a smaller $K$ giving a steeper transition. Therefore, we approximated $g_{\rm{DS}}$ as [^4] Note that here, $g_{\rm{DS}}$ is a function of $T$, not $T_{\rm{DS}}$. This choice was made so Eq. 30 can be used in Eq. 29 to put $T_{\rm{DS}}$ in terms of $T$.
@@
where $T_{\rm{nr}}$ is the visible sector temperature at which $\chi$ becomes non relativistic [^5] We approximated this temperature by using Eq. 29 when all of $\chi$ is non relativistic ($g_{\rm{DS}}(T_{\rm{DS}})=1$). ($m_{\chi}/T_{\rm{DS}}=1$) and we used $g_{\rm{DS},i}=2$ and $g_{\rm{DS},f}=1$ in the following analysis. For each Dark WIMP considered, we determined the steepness of the step function (the value of $K$) using two methods. The first method (columns 2 & 4 in Tab. 2) chose $K$ such that 80% of $\chi$ is non relativistic by $T_{\rm{DS}}\approx\frac{1}{6}m_{\chi}$ [20]. A consequence of this is that $g_{\rm{DS},f}=1$ is not quite reached by the universe today. Our second method for determining $K$ (columns 3 & 5 in Tab. 2) chose $K$ such that $g_{\rm{DS},f}=1$ is reached by $m_{\chi}/T_{\rm{DS}}\approx 10$. It is important to note that the different $g_{\rm{DS}}(T_{\rm{DS}})$ functions used only affect the relic abundance of $\chi$ slightly, and do not have an impact on the phase transition parameters which are entirely determined by $m,\mu$, and $\lambda$. The benchmark cases considered are shown in Tab. 2. The interaction strength ($y^{\prime}$) has been adjusted so the correct relic abundance is achieved and we note that a smaller $K$ results in slightly cooler freeze out temperature ($T_{\rm{FO}}$).
@@
$\langle\sigma_{\chi\chi\rightarrow\xi\xi}v\rangle$ [$\rm{cm^{3}s^{-1}}$] $1.3\times 10^{-26}$ $1.5\times 10^{-26}$ $1.3\times 10^{-27}$ $1.8\times 10^{-27}$
$T_{\rm{FO}}$ [keV] 2.37 2.16 62.78 58.60
@@
Table 2: Benchmark Dark Big Bangs producing Dark WIMP (DW) dark matter with input parameters from Region 2. The input parameters were chosen to feature phase transitions that occur at different times in the universe, are strong enough to fall within the sensitivity bounds of upcoming GW surveys (Fig. 10), and produce the correct relic abundance assuming Dark WIMPs make up all the dark matter in the universe. DW 1 and DW 2 have been split into two columns corresponding to the steepness $K$ of the $g_{\rm{DS}}$ function used (Eq. 30). Benchmark cases featuring the same Dark WIMP masses with parameters chosen from Region 1 can be found in Ref. [1].
The evolution of $\rho_{\chi}$ for each benchmark case is shown in Fig. 9. Once $Y_{\chi}$ is solved for numerically, the dark WIMP energy density is given by $\rho_{\chi}=m_{\chi}Y_{\chi}s$ in the non relativistic regime. The dark WIMP energy density evolution while the particles are in thermal equilibrium and (for the most part) relativistic is given by:
@@
where $K_{1}$ is a modified Bessel function of the second kind and $s$ is the entropy density of the visible sector $s=\frac{2\pi^{2}}{45}g_{\rm{s}}(T)T^{3}$. Eq. 31 is found by numerically integrating the Maxwell-Boltzmann phase space distributions for $n_{\chi,\rm{eq}}$ and $\rho_{\chi,\rm{eq}}$. [^6] As in Ref. [1], we have approximated the Bose-Einstein distribution with a Maxwell-Boltzmann distribution.
Figure 9: Evolution of the Dark WIMP energy density for the benchmark cases shown in Tab. 2. In all cases, the dark matter is produced relativistically and redshifts like $\rho_{\chi}\propto T^{4}$. The Dark WIMPs freeze out as cold relics and join the standard $\Lambda$CDM evolution track, redshifting like $\rho_{\chi}\propto T^{3}$, to produce the correct relic abundance of dark matter in the universe. Constructing $g_{\rm{DS}}(T)$ such that $g_{DS}=1$ by $m_{\chi}/T_{\rm{DS}}\approx 10$ (the $K$ = 0.001, 0.040 dashed lines) results in a small bump in the energy density before Boltzmann suppression sets in, which has been highlighted in the zoomed regions. In both cases, the correct relic abundance is obtained by slightly increasing the interaction strength. The bump also causes an earlier $T_{\rm{FO}}$ (see Tab. 2).
@@
at the time of the DBB. The energy density of gravity waves will subsequently redshift like $\rho_{\rm{GW}}\sim a^{-4}$ [1]. The projected sensitivity windows of the International Pulsar Timing Array (IPTA) [34] and the Square Kilometer Array (SKA) [35] make strong and late DBB’s favorable for detection [1]. Both benchmark Dark Big Bangs fall within IPTA and SKA sensitivity windows, as shown in Fig. 10.
Figure 10: Benchmark DBB phase transition parameters on the projected sensitivity range (colored regions) of the upcoming IPTA and SKA pulsar timing arrays [1]. The hatched regions correspond to the bounds on the DBB discussed in this paper and shown in Fig. 2. BP 1, BP 2, DZ, and DC, are the benchmarks considered in Ref. [1]. Here we have used the $\alpha$ and $T_{*}$ values published in their tables, ignoring the discrepancies mentioned in Fig. 4 since the values tabulated can be achieved with slightly different choices for $m$ and $\mu$. To the figure we have added our new Dark WIMP benchmarks from Region 2: DW 1 and DW 2.
VI Conclusion and Future Directions